# Extending wavelet regularity beyond Gevrey classes

### v1

@@ Abstract
We construct a smooth orthonormal wavelet $\psi$ such that both $\psi$ and its Fourier transform $\widehat{\psi}$ belong to the extended Gevrey class $\mathcal{E}_{\sigma}(\mathbb{R})$ for $\sigma>1$, providing an example that lies beyond all classical Gevrey classes. Our approach uses the idea of invariant cycles to extend the initial Lemarié-Meyer support of the low-pass filter $m_{0}$ from $\frac{2\pi}{3},\frac{2\pi}{3}]$ to $\frac{4\pi}{5},\frac{4\pi}{5}]$. This extension allows us to control the decay rate of $m_{0}$ near $\frac{2\pi}{3}$, which yields global decay estimates for $\psi$ and $\hat{\psi}$. In addition, the decay rates are described using special functions involving the Lambert W function, which plays an important role in our construction.
@@ 1 Introduction
One of the fundamental concepts in wavelet theory is the notion of a multiresolution analysis (MRA in short), introduced by Y. Meyer [17] and later developed by S. Mallat [16]. MRA is based on a hierarchical family of closed subspaces of $L^{2}(\mathbb{R})$ and involves a scaling function $\varphi$, which plays a central role in wavelet construction. Under certain technical assumptions, $\hat{\varphi}$ can be represented as an infinite product constructed from a $2\pi$-periodic function, known as the low-pass filter. More precisely, if $\varphi$ is a scaling function and $m_{0}$ the associated low-pass filter, then
@@
With this approach, it is therefore sufficient to construct a low-pass filter with the desired regularity properties, which are inherited by $\hat{\varphi}$ and $\hat{\psi}$ through (1) and (2). The regularity of $\varphi$ and $\psi$ is then measured via Paley–Wiener type theorems, which relate the smoothness of a function to the decay of its Fourier transform. For a more detailed exposition of the general theory of wavelets and their applications, we refer to [5, 9].
We will use invariant cycles $\left\{-\frac{2\pi}{3},\frac{2\pi}{3}\right\}$ and $\left\{-\frac{2\pi}{5},\frac{2\pi}{5},-\frac{4\pi}{5},\frac{4\pi}{5}\right\}$ to extend the support of the low-pass filter $m_{0}$ on the interval $\pi,\pi)$. For this purpose, let us recall the definition (see [10]).
@@
Let $\rho:[-\pi,\pi)\to[-\pi,\pi):\xi\mapsto 2\xi\ (mod\,2\pi$). For $\xi\in[-\pi,\pi)$, if there exists a number $l\in\mathbb{N}$ such that $\rho^{l}(\xi)=\xi$, we call the corresponding orbit $\mathcal{O}(\xi)=\{\xi,\rho(\xi),\dots,\rho^{l-1}(\xi)\}$ an invariant cycle of length $l$, generated by $\xi$.
The notion of invariant cycles was introduced by A. Cohen to characterize orthonormality conditions for wavelets (see [5]). In [10], it is shown that for a band-limited MRA wavelet, $\prod_{\xi\in\mathcal{O}(\xi)}m_{0}(\xi)=0$ for every invariant cycle $\mathcal{O}(\xi)$. The authors also constructed a smooth low-pass filter with $m_{0}(\pm 2\pi/3)\neq 0$, yielding a non-band-limited wavelet regular only up to finite order. In contrast, Lemarié–Meyer wavelets use smooth low-pass filter that vanish in a neighborhood of $\pm 2\pi/3$, making the corresponding scaling function $\varphi$ and wavelet $\psi$ band-limited and thus analytic in time.
It is well known that a smooth orthonormal wavelet cannot have exponential decay (see [6]). Therefore, constructing a smooth wavelet close to the “critical case”—that is, one which is slightly more regular than $C^{\infty}$, or whose decay is slightly slower than exponential—appears to be a challenging task. An example of such a wavelet can be found in [7], where the authors employ Gevrey regularity and use the translates of the Gevrey function $e^{-|\xi|^{-\frac{1}{s-1}}}$, $s>1$, to control the decay rate of the low-pass filter near $\frac{2\pi}{3}$.
We apply ideas from the cited work to the function
$\begin{split}&f_{\rho,\sigma}(x)=e^{-\rho\,g_{\sigma}\left(1/x\right)},\qquad\rho>0,\;\sigma>1,\\ &\text{where}\\ &g_{\sigma}(x)=\omega_{\sigma}\!\left(\ln\left(1+|x|\right)\right),\qquad x\in\mathbb{R},\\ &\omega_{\sigma}(x)=x^{\frac{\sigma}{\sigma-1}}\big/W^{\frac{1}{\sigma-1}}(x),\qquad x\geq 0,\;\sigma>1,\qquad\omega_{\sigma}(0)=0,\end{split}$ (3)
and $W$ denotes the principal branch of the Lambert $W$ function. This approach yields a new non-band-limited wavelet whose decay rate is faster than polynomial but slower than subexponential. Moreover, its regularity is captured by the extended Gevrey classes.
Extended Gevrey classes of locally smooth functions were introduced in [20]. They contain the union of all Gevrey classes and therefore describe regularity intermediate between Gevrey and $C^{\infty}$. The growth of the derivatives of their elements is controlled by the sequence $\{M^{\tau,\sigma}_{p}\}_{p\in\mathbb{N}_{0}}$, where
@@
for any choice of $\tau>0$ and $\sigma>1$, they cannot be treated within the classical framework of ultradifferentiable function theory. Nevertheless, it has recently been shown that extended Gevrey classes form an important example of weight matrix classes, where regularity is governed by a family of sequences rather than a fixed one (see [11, 23]). Applications of the theory can be found in [3, 12, 13, 2], where extended Gevrey classes are often referred as PTT-spaces (Pilipović, Teofanov, Tomić spaces).
The main purpose of this paper is to establish new regularity properties of wavelets that go beyond the classical theory of ultradifferentiable functions. We refine the result of [24] on band-limited wavelets in extended Gevrey settings by controlling the decay and smoothness of the scaling function in the Fourier domain. At the same time, we obtain a wavelet that is less regular than the one constructed in [7, Theorem 4.4]; equivalently, its regularity lies closer to $C^{\infty}$ (see Lemma 3.6).
Paper is organized as follows: In Subsection 1.1 we fix the notation and discuss the main properties of the Lambert $W$ function. In Section 2, we introduce the extended Gevrey classes $\mathcal{E}_{\sigma}(\mathbb{R})$, discuss their defining sequences $M^{\tau,\sigma}_{p}$ and corresponding associated functions. In Section 2.1 we prove that functions $f_{\rho,\sigma}$ given in (3) belong to $\mathcal{E}_{\sigma}(\mathbb{R})$. Finally, in Section 3 we construct the desired wavelet from the low-pass filter $m_{0}$, by using $f_{\rho,\sigma}$, and prove Theorem 1.2 as our main result:
Theorem 1.2.
Let $\sigma>1$. For $\eta>1$ set $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}\in(1,\sigma_{0})$ where $\sigma_{0}=\max\{2,\sigma\}$. Then there exists an orthonormal wavelet $\psi$ such that
$\psi\in\mathcal{E}_{\sigma}(\mathbb{R})\backslash\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}\mathcal{E}_{\sigma^{\prime}}(\mathbb{R})\quad{\rm and}\quad\displaystyle\hat{\psi}\in\mathcal{E}_{\sigma}(\mathbb{R})\backslash\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}(\mathbb{R}).$
The Appendix contains the proof of Lemma 2.17 and illustrations of the wavelet constructed in Section 3.
@@ 1.1 Preliminaries
Throughout the paper we use the following notation: $\mathbb{N}$, $\mathbb{N}_{0}$, $\mathbb{R}_{+}$, $\mathbb{R}$ and $\mathbb{C}$ denote the sets of natural numbers, non-negative integers, positive real numbers, real numbers, and complex numbers, respectively. We write $f\asymp g$ to denote that the two functions are asymptotically equivalent, meaning that $f=O(g)$ and $g=O(f)$ as $x\to\infty$ (here $f=O(g)$, $x\to\infty$, means that $f(x)\leq L\,(g(x)+1)$ for some $L\geq 1$ and all $x\geq 0$). By $f\prec g$ we mean that $f=o(g)$ as $x\to\infty$, that is, $f(x)/g(x)\to 0$ as $x\to\infty$.
With ${\rm coz}\,f=\{x_{0}\in\mathbb{R}\,|\,f(x_{0})\not=0\}$ we denote a cozero set (complement of a zero set) of the continuous function $f$. Then the support of $f$ is ${\rm supp}\,f=\overline{{\rm\operatorname{coz}}\,f}$, where $\overline{X}$ denotes the closure of the set $X$. The interior of the set $X$ is denoted by ${\rm int}\,X$. We write $\mathbb{K}\subset\subset U$ when $\mathbb{K}$ is an compact subset of an open set $U$.
Let $\nu$ be a non-negative, continuous, increasing and even function on $\mathbb{R}$, with $\nu(0)=0$. Recall (see [1]), $\nu$ is of Braun-Meise-Taylor type (BMT in short) if the following conditions hold:
@@
for $x\in\mathbb{R}$ where $\ln_{+}{|x|}=\max\{0,\ln|x|\}$. Moreover, it was shown in [23] that $\omega_{\sigma}(\ln_{+}|x|)$, $x\in\mathbb{R}$, where $\omega_{\sigma}$ is given in (3), is asymptotically equivalent to a BMT function. Clearly, the same conclusion holds for $g_{\sigma}(x)=\omega_{\sigma}(\ln(1+|x|))$.
We recall some of the basic properties of the Lambert $W$ function. It is defined as the multivalued inverse of the function $z\mapsto ze^{z}$ for $z\in\mathbb{C}$. We will denote its principal branch by $W(z)$, $z\in\mathbb{C}\setminus(-\infty,-e^{-1}]$ (instead of common $W_{0}(z)$). Lambert $W$ function splits the complex plane into infinitely many regions. The boundary curve of its principal branch is given by:
@@
$W(x)\sim x,\quad x\to 0,$ (6)
and $(W4)$ implies that
$W^{\prime}(x)=\frac{W(x)}{x(1+W(x))},\quad x>0.$ (7)
@@ 2 Regularity classes
We begin with a Lemma from [20] that summarizes the main properties of the sequences $M^{\tau,\sigma}_{p}$, $p\in\mathbb{N}_{0},$ given in (4).
@@
Again for $\sigma=1$ we obtain some of the well-known classes. For instance, $\mathcal{E}_{1,1}(\mathbb{R})=\mathcal{A}(\mathbb{R})$, and $\mathcal{E}_{\tau,1}(\mathbb{R})=\mathcal{G}_{\tau}(\mathbb{R})$ for $\tau>1$; these are the spaces of locally analytic functions and Gevrey functions of order $\tau$, respectively.
@@
Next we define a function associated to $M^{\tau,\sigma}_{p}$, $p\in\mathbb{N}_{0}$.
@@
$\displaystyle T_{\tau,\sigma}(x)=\sup_{p\in\mathbf{N}_{0}}\ln\frac{x^{p}}{M^{\tau,\sigma}_{p}},\quad\;\;x>0.$
@@
$\left(M^{\tau,\sigma}_{p}\right)^{1/p}=p^{\tau p^{\sigma-1}}$
is clearly bounded below by a positive constant for any $\tau>0$ and $\sigma>1$, it follows that $T_{\tau,\sigma}(x)$ vanishes for sufficiently small $x>0$ (see [14]). Therefore we may assume $T_{\tau,\sigma}(0)=0$.
It turns out that for any fixed $\tau>0$, $T_{\tau,\sigma}(|x|)$ is asymptotically equivalent to the function $g_{\sigma}(x)=\frac{\ln^{\frac{\sigma}{\sigma-1}}(1+x)}{W^{\frac{1}{\sigma-1}}(\ln(1+x))}$ which appears in (3). In particular, following Proposition holds (see [22, 23]).
@@
The following definition introduces the decay rate studied in this paper.
@@
Let $\sigma>1$ and $g_{\sigma}$ be as in (3). A continuous function $\phi$ belongs to ${{\mathbf{\Gamma}}}_{\sigma}\left(\mathbb{R}\right)$ if
$(\exists\rho>0)\,(\exists C>0)\quad\quad|\phi(\xi)|\leq C\,e^{-\rho g_{\sigma}(\xi)},\quad\xi\in\mathbb{R}.$ (12)
We can now formulate the Paley–Wiener theorem, which relates the regularity of $\mathcal{E}_{\sigma}$ to the decay rate given by ${{\mathbf{\Gamma}}}_{\sigma}$. For further details, we refer to [22, 25].
@@
Remark 2.12.
Part $a)$ of Theorem 2.11 can be easily extended to smooth functions that are not compactly supported but have sufficient decay at infinity. In particular, if a smooth function $\varphi$ satisfies $\widehat{\varphi}\notin\mathbf{\Gamma}_{\sigma}$, then $\varphi\notin\mathcal{E}_{\sigma}$.
We conclude this section with the following technical lemma, which will be used later.
@@
$\leq C_{1}e^{-\frac{\rho}{2}\,g_{\sigma}(\xi)},\quad\xi\geq\frac{1}{2},$
@@
which follows easily by (11). The other implication can be proved in a similiar way. ∎
@@
$M^{\tau,\sigma}_{p}=p^{\tau p^{\sigma}},\qquad p\in\mathbb{N}_{0},$
satisfies a strong non-quasianalyticity condition for every choice of $\tau>0$ and $\sigma>1$:
@@
Moreover, if $a^{\tau,\sigma}_{p}=\frac{M^{\tau,\sigma}_{p+1}}{M^{\tau,\sigma}_{p}}$, they proved the stronger assertion that the gamma index of the sequence $\mathbb{M}^{\tau,\sigma}=\{M^{\tau,\sigma}_{p}\}_{p\in\mathbb{N}_{0}}$,
$\gamma(\mathbb{M}^{\tau,\sigma})=\sup\left\{\gamma>0:\;(\exists C>0)(\forall p\in\mathbb{N}_{0})(\forall q\geq p)\,\frac{a^{\tau,\sigma}_{p}}{(p+1)^{\gamma}}\leq C\,\frac{a^{\tau,\sigma}_{q}}{(q+1)^{\gamma}}\right\},$
@@
This implies that the corresponding function classes admit (ultraholomorphic) extensions to unbounded sectors in $\mathbb{C}$ of the form
@@
for arbitrary $\gamma>0$, where $\mathcal{R}$ denotes the Riemann surface of the logarithm. Such extensions will be used in Section 2.1 to construct an example in $\mathcal{E}_{\sigma}(\mathbb{R})$.

### v2

@@ Abstract
We present a general construction of smooth orthonormal wavelet $\psi$ which, together with its Fourier transform $\widehat{\psi}$ belongs to the extended Gevrey class $\mathcal{E}_{\sigma}(\mathbb{R})$ for $\sigma>1$, providing an example that lies beyond all classical Gevrey classes. Our approach uses the idea of invariant cycles to extend the initial Lemarié-Meyer support of the low-pass filter $m_{0}$ from $\frac{2\pi}{3},\frac{2\pi}{3}]$ to $\frac{4\pi}{5},\frac{4\pi}{5}]$. This extension allows precise control of the the decay rate of $m_{0}$ near $\frac{2\pi}{3}$, which yields global decay estimates for $\psi$ and $\hat{\psi}$. In addition, the decay rates are described using special functions involving the Lambert $W$ function, which plays an important role in our construction.
@@ 1 Introduction
In this paper, we present a construction of smooth, non-band-limited orthonormal wavelets $\psi$ whose regularity is weaker than Gevrey regularity both in time and in frequency. This is achieved by controlling the decay of the low-pass filter $m_{0}$ near the invariant cycle points $\pm 2\pi/3$ [8]. For this purpose, we use certain flat functions and their extensions to complex domain, see [12, 13, 14]. Our methods differ from those used in [26] where band-limited wavelets were constructed instead.
The notion of a multiresolution analysis (MRA) is one of the fundamental concepts in wavelet theory, introduced by Y. Meyer [18] and later developed by S. Mallat [17]. MRA is based on a hierarchical family of closed subspaces of $L^{2}(\mathbb{R})$ and involves a scaling function $\varphi$, which plays a central role in construction of the orthonormal wavelet bases. Under certain technical assumptions, its Fourier transform $\hat{\varphi}$ can be represented as an infinite product constructed from a $2\pi$-periodic function, known as the low-pass filter. More precisely, if $\varphi$ is a scaling function and $m_{0}$ the associated low-pass filter, then
@@
In such framework, it is therefore sufficient to construct a low-pass filter with the desired regularity properties, which are then inherited by $\hat{\varphi}$ and $\hat{\psi}$ through (1) and (2). The regularity of $\varphi$ and $\psi$ is then given by Paley–Wiener type theorems, which relate the smoothness of a function to the decay of its Fourier transform. For a more detailed exposition of the general theory of wavelets and their applications, we refer to classical textbooks [5, 10, 19].
We shall use invariant cycles $\left\{-\frac{2\pi}{3},\frac{2\pi}{3}\right\}$ and $\left\{-\frac{2\pi}{5},\frac{2\pi}{5},-\frac{4\pi}{5},\frac{4\pi}{5}\right\}$ to extend the support of the low-pass filter $m_{0}$ on the interval $\pi,\pi)$. For this purpose, let us recall the definition of an invariant cycle (see [11]).
@@
Let $\rho:[-\pi,\pi)\to[-\pi,\pi):\xi\mapsto 2\xi\ (mod\,2\pi$). For $\xi\in[-\pi,\pi)$, if there exists $l\in\mathbb{N}$ such that $\rho^{l}(\xi)=\xi$, we call the corresponding orbit $\mathcal{O}(\xi)=\{\xi,\rho(\xi),\dots,\rho^{l-1}(\xi)\}$ an invariant cycle of length $l$, generated by $\xi$.
The notion of invariant cycles was introduced by A. Cohen to characterize orthonormality conditions for wavelets (see [5]). In [11], it is proved that for a band-limited MRA wavelet, $\prod_{\xi\in\mathcal{O}(\xi)}m_{0}(\xi)=0$ for every invariant cycle $\mathcal{O}(\xi)$, and furthermore the authors constructed a smooth low-pass filter with $m_{0}(\pm 2\pi/3)\neq 0$, yielding a non-band-limited wavelet regular only up to a finite order. In contrast, Lemarié–Meyer (see [10]) use smooth low-pass filter that vanishes in a neighborhood of $\pm 2\pi/3$, making the corresponding scaling function $\varphi$ and wavelet $\psi$ band-limited and thus analytic in time. In our construction, $m_{0}(\pm 2\pi/3)=0$ and $m_{0}(\xi)\neq 0$ in a neighborhood of $\pm 2\pi/3$ ($\xi\neq\pm 2\pi/3$), yielding to non-band-limited scaling function $\varphi$ and wavelet $\psi$ with extended Gevrey regularity.
It is well known that a smooth orthonormal wavelet cannot have exponential decay (see [6]). Therefore, constructing a smooth wavelet whose decay is slower than exponential appears to be a challenging task (see [7, 21]). An example of such a wavelet can be found in [8], where the authors employ Gevrey regularity and use the translates of the Gevrey function $e^{-|\xi|^{-\frac{1}{s-1}}}$, $s>1$, to control the decay rate of the low-pass filter near $\frac{2\pi}{3}$.
We employ the ideas from [8] and consider functions of the form
$\begin{split}&f_{\rho,\sigma}(x)=e^{-\rho\,g_{\sigma}\left(1/x\right)},\qquad\rho>0,\;\sigma>1,\;x\not=0,\\ &\text{where}\\ &g_{\sigma}(x)=\omega_{\sigma}\!\left(\ln\left(1+|x|\right)\right),\\ &\omega_{\sigma}(x)=x^{\frac{\sigma}{\sigma-1}}\big/W^{\frac{1}{\sigma-1}}(x),\quad\omega_{\sigma}(0)=0,\end{split}$ (3)
where $W$ denotes the principal branch of the Lambert $W$ function (see subsection 1.1). This approach yields a new class of non-band-limited orthonormal wavelets whose decay rate is faster than polynomial but slower than subexponential. Moreover, its regularity in time and frequency domains is captured by the extended Gevrey classes.
Extended Gevrey classes of locally smooth functions were introduced in [22]. They contain the union of all Gevrey classes and therefore describe an intermediate regularity between Gevrey and $C^{\infty}$. The growth of the derivatives of their elements is controlled by the sequence $\{M^{\tau,\sigma}_{p}\}_{p\in\mathbb{N}_{0}}$, where
@@
for any choice of $\tau>0$ and $\sigma>1$, the extended Gevrey classes cannot be treated within the classical framework of ultradifferentiable function theory. Nevertheless, it has recently been shown that extended Gevrey classes form an important example of weight matrix classes, where regularity is governed by a family of defining sequences (see [12, 25]). Applications of the theory can be found in [3, 13, 14, 2], where extended Gevrey classes are referred to as PTT-spaces.
The main purpose of this paper is to establish new regularity properties of orthonormal wavelets that go beyond the classical theory of ultradifferentiable functions. We refine the result of [26] on band-limited wavelets in extended Gevrey settings by controlling the decay and smoothness of the scaling function in the Fourier domain. At the same time, we obtain a wavelet that is less regular than the one constructed in [8, Theorem 4.4]; equivalently, its regularity lies closer to $C^{\infty}$ (see Lemma 3.7).
Paper is organized as follows: In Subsection 1.1 we fix the notation and discuss the main properties of the Lambert $W$ function. In Section 2, we introduce the extended Gevrey classes $\mathcal{E}_{\sigma}(\mathbb{R})$, see Definition 2.3, discuss their defining sequences $M^{\tau,\sigma}_{p}$ and corresponding associated functions. In particular, we prove that functions $f_{\rho,\sigma}$ given in (3) belong to $\mathcal{E}_{\sigma}(\mathbb{R})$, see Theorem 2.17 and Proposition 2.18. In Section 3 we construct the desired wavelet from the low-pass filter $m_{0}$, by using $f_{\rho,\sigma}$, and prove Theorem 3.5 as our main result. This implies the following particular result:
Theorem.
There exists an orthonormal wavelet $\psi$ such that for arbitrary $\eta>1$
$\psi\in\mathcal{E}_{2}(\mathbb{R})\backslash\bigcup_{1<\sigma^{\prime}<\frac{2+\eta}{1+\eta}}\mathcal{E}_{\sigma^{\prime}}(\mathbb{R})\quad{\rm and}\quad\displaystyle\hat{\psi}\in\mathcal{E}_{2}(\mathbb{R})\backslash\bigcup_{1<\sigma^{\prime}<2}\mathcal{E}_{\sigma^{\prime}}(\mathbb{R}).$
To illustrate our construction, we present several graphs of ONW given by Theorem 3.5 in Subsection 3.1.
@@ 1.1 Preliminaries
Throughout the paper we use the following notation: $\mathbb{N}$, $\mathbb{N}_{0}$, $\mathbb{R}_{+}$, $\mathbb{R}$ and $\mathbb{C}$ denote the sets of natural numbers, non-negative integers, positive real numbers, real numbers, and complex numbers, respectively. We write $f\asymp g$ to denote that the two functions are asymptotically equivalent, meaning that $f=O(g)$ and $g=O(f)$ as $x\to\infty$ (here $f=O(g)$, $x\to\infty$, means that $f(x)\leq L\,(g(x)+1)$ for some $L\geq 1$ and all $x\geq 0$). By $f\prec g$ we mean that $f=o(g)$ as $x\to\infty$, that is, $f(x)/g(x)\to 0$ as $x\to\infty$. We also write $f\sim g$ when $f(x)/g(x)\to 1$ as $x\to\infty$.
With ${\rm coz}\,f=\{x_{0}\in\mathbb{R}\,|\,f(x_{0})\not=0\}$ we denote a cozero set (complement of a zero set) of the continuous function $f$. Then the support of $f$ is ${\rm supp}\,f=\overline{{\rm\operatorname{coz}}\,f}$, where $\overline{X}$ denotes the closure of the set $X$. The interior of the set $X$ is denoted by ${\rm int}\,X$. We write $\mathbb{K}\subset\subset U$ when $\mathbb{K}$ is a compact subset of an open set $U$.
Let $\nu$ be a non-negative, continuous, increasing and even function on $\mathbb{R}$, with $\nu(0)=0$. Then (see [1]), $\nu$ is of Braun-Meise-Taylor type (BMT in short) if the following conditions hold:
@@
for $x\in\mathbb{R}$ where $\ln_{+}{|x|}=\max\{0,\ln|x|\}$. Moreover, it was shown in [25, Theorem 1] that $\omega_{\sigma}(\ln_{+}|x|)$, $x\in\mathbb{R}$, where $\omega_{\sigma}$ is given in (3), is asymptotically equivalent to a BMT function.
Next, we recall some of the basic properties of the Lambert $W$ function. It is defined as the multivalued inverse of the function $z\mapsto ze^{z}$ for $z\in\mathbb{C}$. We will denote its principal branch by $W(z)$, $z\in\mathbb{C}\setminus(-\infty,-e^{-1}]$, which is also denoted by $W_{0}(z)$ in the literature. Lambert $W$ function splits the complex plane into infinitely many regions. The boundary curve of its principal branch is given by:
@@
$W(x)\sim x,\quad x\to 0.$ (7)
@@ 2 Regularity classes
We begin with a lemma that summarizes the main properties of the sequences $M^{\tau,\sigma}_{p}$, $p\in\mathbb{N}_{0},$ given in (4). We refer to [22] for the proof, see also [27]
@@
Again for $\sigma=1$ we obtain some of the well-known classes. For instance, $\mathcal{E}_{1,1}(\mathbb{R})=\mathcal{A}(\mathbb{R})$, and $\mathcal{E}_{\tau,1}(\mathbb{R})=\mathcal{G}_{\tau}(\mathbb{R})$ for $\tau>1$ are the spaces of locally analytic functions and Gevrey functions of order $\tau$, respectively. Recall, $\phi\in\mathcal{G}_{\tau}(\mathbb{R})$ if
$(\forall\,\mathbb{K}\subset\subset\mathbb{R})(\exists C>0)\quad\sup_{x\in\mathbb{K}}|\phi^{(p)}(x)|\leq C^{p+1}p!^{\tau},\,\,p\in\mathbb{N}_{0}.$
@@
Next we define a function associated to the sequence $M^{\tau,\sigma}_{p}$, $p\in\mathbb{N}_{0}$.
@@
$\displaystyle T_{\tau,\sigma}(x)=\sup_{p\in\mathbb{N}_{0}}\ln\frac{x^{p}}{M^{\tau,\sigma}_{p}},\quad\;\;x>0.$
@@
$\left(M^{\tau,\sigma}_{p}\right)^{1/p}=p^{\tau p^{\sigma-1}},\quad p\in\mathbb{N},$
is clearly bounded below by a positive constant for any $\tau>0$ and $\sigma>1$, it follows that $T_{\tau,\sigma}(x)$ vanishes for sufficiently small $x>0$ (see [15]). Thus we may consider $T_{\tau,\sigma}(|x|)$, $x\in\mathbb{R}$, instead.
It turns out that for any fixed $\tau>0$, $T_{\tau,\sigma}(|x|)$ is asymptotically equivalent to the function $g_{\sigma}(x)=\frac{\ln^{\frac{\sigma}{\sigma-1}}(1+|x|)}{W^{\frac{1}{\sigma-1}}(\ln(1+|x|))}$ which appears in (3). In particular, following Proposition holds (see [24, 25]).
@@
The following definition introduces the precise decay rate studied in this paper.
@@
Let $\sigma>1$ and $g_{\sigma}(x)=\frac{\ln^{\frac{\sigma}{\sigma-1}}(1+|x|)}{W^{\frac{1}{\sigma-1}}(\ln(1+|x|))}$, $x\in\mathbb{R}$. A continuous function $\phi$ belongs to ${{\mathbf{\Gamma}}}_{\sigma}\left(\mathbb{R}\right)$ if
$(\exists\rho>0)\,(\exists C>0)\quad\quad|\phi(x)|\leq C\,e^{-\rho g_{\sigma}(x)},\quad x\in\mathbb{R}.$ (12)
We can now state the Paley–Wiener theorem, which characterizes the regularity of $\mathcal{E}_{\sigma}$ in terms of the decay rate given by $\mathbf{\Gamma}_{\sigma}$. For further details, we refer to [24, 27].
@@
We conclude this section with the following technical lemma, which will be used in the sequel.
@@
$\leq C_{1}e^{-\frac{\rho}{2}\,g_{\sigma}(\xi)},\quad\quad|\xi|\geq 1,$
@@
which follows easily by (11). The other implication is trivial. ∎
@@
$M^{\tau,\sigma}_{0}=1\quad\text{and}\quad M^{\tau,\sigma}_{p}=p^{\tau p^{\sigma}},\qquad p\in\mathbb{N},$
satisfies a strong non-quasianalyticity condition for every choice of $\tau>0$ and $\sigma>1$, i.e.,
@@
Moreover, if $m^{\tau,\sigma}_{p}=\frac{M^{\tau,\sigma}_{p+1}}{M^{\tau,\sigma}_{p}}$, $p\in\mathbb{N}_{0}$, the stronger assertion holds: the gamma index of the sequence $\mathbb{M}^{\tau,\sigma}=\{M^{\tau,\sigma}_{p}\}_{p\in\mathbb{N}_{0}}$, given by
$\gamma(\mathbb{M}^{\tau,\sigma})=\sup\left\{\gamma>0:\;(\exists C>0)(\forall p\in\mathbb{N}_{0})(\forall q\geq p)\,\frac{m^{\tau,\sigma}_{p}}{(p+1)^{\gamma}}\leq C\,\frac{m^{\tau,\sigma}_{q}}{(q+1)^{\gamma}}\right\},$
@@
This implies that the corresponding function classes $\mathcal{E}_{\tau,\sigma}(\mathbb{R})$, $\tau>0$, $\sigma>1$, admit (ultraholomorphic) extensions to unbounded sectors in $\mathbb{C}$ of the form
@@
for arbitrary $\gamma>0$, where $\mathcal{R}$ denotes the Riemann surface of the logarithm.